# Exact gradients for linear optics with single photons

### v1

@@ III.3 Generalized parameter-shift rule
$f^{\prime}(\theta)=\sum_{\mu=1}^{2n}f(\theta+\theta_{\mu})\dfrac{(-1)^{\mu+1}}{4R\sin^{2}(\theta_{\mu}/2)}\,,$ (27)
@@
A proof of the Corollary is provided in Appendix B.1.
@@ V Applications
In this section, we demonstrate two applications where the GPSR is employed and compared against alternative approaches to computing gradients. In the first example, we focus on a QML task that can be run in a linear optical setup, and compare using the GPSR against FD methods to calculate gradients. We show that using the GPSR can improve training compared to FD methods. In the second example, we instead show how the GPSR could be used to accurately study a Bell state generation circuit.
@@ V.1 QML classification
To train our circuit parameters, we need to calculate the gradient of the loss function with respect to the $\theta_{k}$ parameters. Hence, we need to compute gradients of the function $f$ with respect to beamsplitter parameters. To achieve this we use either the GPSR or FD, with the same total number of samples $N_{\text{tot}}$ (600 per calculation). For a single parameter update step, stochastic gradient descent is used to optimize the $\boldsymbol{\theta}$ angles with the $\boldsymbol{\alpha}$ parameters held constant, and then Adam is used to optimize the classical parameters $\boldsymbol{\alpha}$, with $\boldsymbol{\theta}$ held constant.
@@
Accuracy [%] $89.4\pm 0.43$ $79.9\pm 2.4$ $89\pm 1.2$ $88.5\pm 0.7$
Table 1: Binary classification accuracy. We plot the average prediction accuracy for the 10 models trained using each method, and the standard error. The GPSR leads to a slightly better average and lower variance than the FD alternatives.
In Fig. 6, we plot the loss function during training, averaged over ten different random initializations. We observe that the lowest losses throughout training were achieved by models using the GPSR. The FD trained models had higher variance in the loss curves, with a step size of $h=0.1$ being heuristically optimal. Although the losses for these optimal FD models converged to a similar value as those of the GPSR models, we see that convergence occurred faster for those using the GPSR. All models were trained to achieve high classification accuracies (see: Tab. 1). The GPSR did not result in a statistically significant accuracy increase in the mean, however we see that the accuracies achieved with the GPSR trained models were more consistent throughout the 10 different initializations. The model accuracies achieved may be further improved with further refinement of the hyperparameters used, however these results are already sufficient to highlight a practical difference between the GPSR and FD methods. As discussed in Appendix B and observed in Fig. 6, step sizes for FD methods need to be fine-tuned depending on the task considered to achieve accurate gradient calculation; the GPSR will always be optimal. As such, the GPSR stands out from FD methods in that it can be applied to achieve good training results without prior knowledge about the problem.
@@ VI Discussion
We also note that, for optimization tasks, while employing the GPSR such that it is accurate enough might require more samples, there might be a benefit in convergence speed. However, we leave as future work to explore this trade-off in more detail, as well as strategies to try and avoid poor local minima [68, 69] and barren plateaus [70, 71] which were not discussed here. Possible future works may investigate how to combine the GPSR for linear optics with stochastic approaches [40, 48, 49]. A more detailed analysis of the approximation capabilities of Dirichlet kernels could also bring insight on when is it appropriate to truncate the number of Fourier frequencies, and hence parameter shifts.
We also expect that in the presence of photon loss [72, 73, 74, 75], the GPSR will still be valid, and the derivative of the noisy observable may be accurately reconstructed. Indeed, a lost photon will effectively reduce the number of frequencies in the function, and trigonometric interpolation for $n$ frequencies works also for Fourier series with $(n-1)$ frequencies. Similarly, linear optics with partially distinguishable photons [76] was demonstrated to be equivalent to boson sampling with fewer ideal photons [77], so we expect the GPSR to hold as well. Finally, having introduced a way to assess the behaviour of a broad family of observables with respect to parameters in the circuit, we anticipate that this might serve as a resource to better understand the capabilities and limitations of bosonic systems.

### v2

@@ III.3 Generalized parameter-shift rule
$f^{\prime}(\theta)=\sum_{\mu=1}^{2n}f(\theta+\theta_{\mu})\dfrac{(-1)^{\mu+1}}{4n\sin^{2}(\theta_{\mu}/2)}\,,$ (27)
@@
A proof of the Corollary is provided in Appendix B.1. In Appendix A.2, we generalize the main results of both Theorem 1 and Corollary 1 to more realistic, lossy setups. We show that similarly to the ideal scenario, states subject to loss in linear optical systems still allow for a Fourier expansion which contains at most $n$ positive frequencies. Hence, the same results can also be applied to realistic experimental scenarios.
@@ V Applications
In this section, we demonstrate two applications where the GPSR is employed and compared against alternative approaches to computing gradients. In the first example, we focus on a QML task that can be run in a linear optical setup, and compare using the GPSR against FD methods to calculate gradients. We show that using the GPSR can result in more reliable training compared to FD methods. In the second example, we instead show how the GPSR could be used to accurately study a Bell state generation circuit.
@@ V.1 QML classification
To train our circuit parameters, we need to calculate the gradient of the loss function with respect to the $\theta_{k}$ parameters. Hence, we need to compute gradients of the function $f$ with respect to beamsplitter parameters. To achieve this, we use either the GPSR or FD, with the same total number of samples $N_{\text{tot}}$ (600 per gradient calculation). In the case of the GPSR, the samples are allocated according to the 1-norm of the dirichlet kernels. For a single parameter update step, stochastic gradient descent is used to optimize the $\boldsymbol{\theta}$ angles with the $\boldsymbol{\alpha}$ parameters held constant, and then Adam is used to optimize the classical parameters $\boldsymbol{\alpha}$, with $\boldsymbol{\theta}$ held constant.
@@
Accuracy [%] $90.25\pm 0.21$ $83.35\pm 0.93$ $90.14\pm 0.25$ $89.17\pm 0.31$
Table 1: Binary classification accuracy. We tabulate the average prediction accuracy for the 30 models trained using each method, and the standard error. We see that the GPSR and optimal FD trained models achieve the best classification accuracies.
In Fig. 6, we plot the loss function during training, averaged over thirty different random initializations. We observe that the lowest losses throughout training were achieved by models using the GPSR and FD with the heuristically optimal shift of $0.1$. The FD trained models with non-optimal shift size had higher variance in the loss curves and converged to higher loss values. All models were trained to achieve high classification accuracies (see: Tab. 1), and we see that both the GPSR and the optimal FD models achieved the highest accuracies, with no statistically significant differences between the two. The model accuracies achieved may be further improved with further refinement of the hyperparameters used, however these results are already sufficient to highlight a practical difference between the GPSR and FD methods with different shift sizes. As discussed in Appendix B and observed in Fig. 6, step sizes for FD methods need to be fine-tuned depending on the task considered to achieve accurate gradient calculation; the GPSR will always be optimal. As such, the GPSR stands out from FD methods in that it can be applied to achieve good training results without prior knowledge about the problem.
@@ VI Discussion
We also note that, for optimization tasks, while employing the GPSR may require more samples to yield an accurate gradient, there might be a benefit in convergence speed. However, we leave as future work to explore this trade-off in more detail, as well as strategies to try and avoid poor local minima [68, 69] and barren plateaus [70, 71] which were not discussed here. Possible future works may investigate how to combine the GPSR for linear optics with stochastic approaches [40, 48, 49]. A more detailed analysis of the approximation capabilities of Dirichlet kernels could also bring insight on when is it appropriate to truncate the number of Fourier frequencies, and hence parameter shifts.
An important experimental consideration is the presence of imperfections such as photon loss or partial distinguishability. We anticipate that the GPSR is still valid for most common sources of imperfections. For instance, in Appendix A.2 we show that the GPSR protocol holds also for interferometers with photon loss. Similarly, linear optics with partially distinguishable photons [72] was demonstrated to be equivalent to boson sampling with fewer ideal photons [73], so the GPSR will hold as well.
Note added: During the preparation of this manuscript, we became aware of recent work in Ref. [74] in which the authors present a similar result for a restricted version of the GPSR for circuits involving two photons, with a supplementary note deriving a generalization similar to ours. Our work goes further by giving an in-depth analysis of the main result, including sample complexity scaling estimations and protocols to reduce the number of parameter shifts. We also acknowledge concurrent work in Refs. [75, 76].
@@ A.1 First order derivative reconstruction
A.2 Presence of noise in the system
In this section, we discuss the presence of noise in the system. Experimental realizations of linear optical circuits are in fact faced with photon loss in the system. For a system with $n$ photons, the architecture has some probability of losing a certain amount of photons, a probability which generally depends on the specifics of the platform. For uniform loss with transmission coefficient $\eta$, loss commutes with the unitary $\mathcal{U}(\boldsymbol{\theta})$ and therefore can be applied directly to the initial state $\rho_{\boldsymbol{n}}:=\ket{\boldsymbol{n}}\bra{\boldsymbol{n}}$ | [77]. We define the transformation where a photon in a given mode escapes and is injected into the environment
$a^{\dagger}\mapsto\sqrt{\eta}a^{\dagger}+\sqrt{1-\eta}e^{\dagger}$ (68)
where $e^{\dagger}$ is the creation operator of the environment. The action on an $n$-photon state is given by
$\begin{split}\dfrac{1}{\sqrt{n!}}(a^{\dagger})^{n}\ket{0}_{a}\otimes\ket{0}_{e}&\mapsto\dfrac{1}{\sqrt{n!}}\bigg{(}\sqrt{\eta}a^{\dagger}+\sqrt{1-\eta}e^{\dagger}\bigg{)}^{n}\ket{0,0}\\ &=\dfrac{1}{\sqrt{n!}}\sum_{k=0}^{n}\binom{n}{k}\eta^{k/2}(1-\eta)^{\frac{n-k}{2}}\sqrt{k!}\sqrt{(n-k)!}\ket{k,n-k}\\ &=\sum_{k=0}^{n}\binom{n}{k}^{1/2}\eta^{k/2}(1-\eta)^{\frac{n-k}{2}}\ket{k,n-k}\,.\end{split}$ (69)
Tracing out the environment, we get that the pure loss channel acts on an $n$-photon state $\rho_{n}=\ket{n}\bra{n}$ | as
$\Lambda_{\eta}(\rho_{n}):=\sum_{k=0}^{n}\binom{n}{k}\eta^{k}(1-\eta)^{n-k}\rho_{k}=\sum_{k=0}^{n}p^{\eta}_{n,k}\rho_{k}\,,$ (70)
where we deonte by $p^{\eta}_{n,k}$ the probability of being in state $\rho_{k}:=\ket{k}\bra{k}$ |. Now, the same channel acting on $m$ modes gives
$\Lambda_{\eta}^{\otimes m}(\rho_{\boldsymbol{n}})=\bigotimes_{i=1}^{m}\bigg{(}\sum_{k_{i}=0}^{n_{i}}p^{\eta}_{n_{i},k_{i}}\rho_{k_{i}}\bigg{)}=\sum_{k_{1}=0}^{n_{1}}\dots\sum_{k_{m}=0}^{n_{m}}\bigg{(}\prod_{i=1}^{m}p^{\eta}_{n_{i},k_{i}}\bigg{)}\ket{\boldsymbol{k}}\bra{\boldsymbol{k}}=\sum_{\boldsymbol{k}}p^{\eta}_{\boldsymbol{n},\boldsymbol{k}}\ket{\boldsymbol{k}}\bra{\boldsymbol{k}}\,,$ (71)
where $p^{\eta}_{\boldsymbol{n},\boldsymbol{k}}$ denotes the product in the parentheses, and the final sum runs over all |$\ket{\boldsymbol{k}}$ Fock states such that $k_{i}\leq n_{i}$. The expectation value of an observable $O$ measured with a parametrized, noisy device will be
[…]
where the $c_{\omega,k}$’s terms, similarly to the lossless case, depend on the unitary and the observable. In the second equality we extended the sum by simply considering $c_{\omega,k}\equiv 0$ for $|\omega|>k$. The expectation value then takes the form
$\begin{split}\expectationvalue{O}_{\eta}=\sum_{\omega=-n}^{n}\bigg{(}\sum_{k=0}^{n}c_{\omega,k}\binom{n}{k}\eta^{k}(1-\eta)^{n-k}\bigg{)}e^{i\omega\theta}=\sum_{\omega=-n}^{n}c_{\omega,\eta}e^{i\omega\theta}\,.\end{split}$ (74)
This form of the expectation value shows that all the results derived for the ideal case also hold in the case of circuits with uniform loss.
In the case of mode-dependent losses, the dynamics becomes more complex as the losses cannot commute and be shifted to the beginning (or end) of the circuit. However, it is still reasonable to assume that, in practice, we end up with the same Fourier spectrum. In fact, very generally, we can consider the unitary dilation of a lossy linear optics circuit, where the final state of the $m$ modes is given by
[…]
where $E$ denotes the entire environment system, and $\widetilde{\mathcal{U}}(\theta,\boldsymbol{\eta})$ is a unitary specified from the ideal circuit, as well as additional beamsplitters coupling the $m$ modes and the environment (which act in the same way as Eq. (75), with distinct parameters $\eta_{i}$). We denote $\boldsymbol{\eta}=(\eta_{1},...\eta_{m})$ the loss parameters. An observable measured on the $m$ modes yields the following expectation value
[…]
While needing to consider an enlarged space containing environment modes, and more complex dynamics which include the loss events, the state $\rho_{\boldsymbol{n}}\otimes\ket{\boldsymbol{0}}\bra{\boldsymbol{0}}_{E}$ is still a $n$-photon state and the unitary can still be cast so that the parametrized operation can be singled out and expanded to achieve a spectrum $\{-n,\dots,n\}$.